OpenStudy (anonymous):

Which is an equation in point-slope form of the line that passes through the points ( -4, -1) and (5, 7) ? Please help me

OpenStudy (anonymous):

y = mx + b m = (y2 - y1)/ (x2 - x1) solve for m and then use any of these two points to figure out b.

OpenStudy (anonymous):

The options are A. Y+4=9/8 (x+1). B. Y+1=8/9 (x+4) C. Y+1=8/7 (x+4) D. Y-1=8/9 (x-4)

OpenStudy (anonymous):

@Boblovesmath